# SWE-SQL: Illuminating LLM Pathways to Solve User SQL Issues in Real-World Applications

### v3

@@ 3 BIRD-CRITIC Benchmark
BIRD-CRITIC is developed via a multi-stage annotation converting raw user issues into executable, verifiable tasks. This involves two annotation groups: 1) 10 qualified database/SQL annotators, who pass strict entry test as detailed in Appendix 11.1 and systematic training shown in Appendix 11.2 to promise the quality of annotation; 2) 3 senior database experts/scientists for final data collection decision. This process is visually outlined in Figure 2.
@@
We leverage relational databases from the BIRD-SQL development set [26] chosen for its domain diversity across real data-science tasks (California Schools, Financial, Superheroes) and its permissive license. We migrate their original SQLite schemas to PostgreSQL, MySQL, SQL Server, and Oracle, four widely used production-grade dialects. During migration, we go beyond direct dialect translation by refining table and column names. We adjust data types and introduce guarded alterations to schema components to reduce potential information leakage (see Appendix 10.2). To pair these databases with realistic debugging scenarios, we collect SQL issue queries from Stack Overflow, following a strict protocol shown in Appendix 10.3.
@@
Annotators carefully review the reproduced issue $(\mathcal{P},\mathcal{S},\sigma_{\text{issue}})$ and craft a new $\sigma^{*}$. This annotation requires ensuring that $\sigma^{*}$ can accurately fulfill the user’s objective as inferred from $\mathcal{P}$ and the context of $\sigma_{\text{issue}}$. Also, to ensure robust evaluation, each task is annotated with evaluation scripts consisting of specific test cases written by Python and SQLs. Details can be found in Appendix 12. We report the Success Rate (SR %), considering a task solved only when $\sigma_{\text{pred}}$ successfully passes all test cases in its evaluation script.
@@
Figure 3: Distribution of issue categories in all BIRD-CRITIC, derived from an analysis of SQL usage in the real-world database applications. A detailed distribution is in Appendix 14.
@@
Table 3 summarizes the key properties of the BIRD-CRITIC benchmark, and Figure 3 visualizes the distribution of its underlying knowledge categories. The distribution of benchmark, is detailed in Appendix 14. A side-by-side comparison with standard text-to-SQL benchmarks (Table 6 in Appendix 14.1) exposes three distinctive challenges introduced by BIRD-CRITIC: non-query-like problems, multi-dialect complexities, and the most verbose but authentic user queries. As far as we know, BIRD-CRITIC is the first debugging benchmark for SQL applications. These aspects establish BIRD-CRITIC as a crucial benchmark for rigorously evaluating LLM proficiency in solving authentic SQL issues.
@@ 4 Six-Gym: An Automated SQL Debugging Environment for LLMs
GYM-like datasets have proven effective for training LLMs as agents for complex tasks [30]. However, manually collecting and annotating these datasets is labor-intensive and difficult to scale, especially for debugging tasks. Thus, we introduce SQL-Rewind, which addresses this by inverting the debugging paradigm: starting with correct SQL queries ($\sigma^{*}$) and systematically introducing realistic issues to generate issue SQLs ($\sigma_{\text{issue}}$) and user issue query $\mathcal{P}$. This approach enables efficient creation of large-scale training data without human annotation. The pseudo-algorithm is shown in Appendix 17.1.
@@ 6.1 SetUp
We evaluate the performance of several popular and strong LLMs across two primary categories, including general-purpose models: Gemini-2.0-Flash, GPT-4.1, Claude-3.7-Sonnet, Qwen-2.5-Coder-32B, Meta-Llama-3.1-8B, Meta-Llama-3.3-70B, Phi-4 and DeepSeek-V3. The second category consists of models specifically renowned for their advanced reasoning capabilities: O3-mini, O1-preview, Gemini-2.0-Flash-Thinking, and Claude-3.7-Sonnet-Thinking. The implementation details are in Appendix 16.2.
@@ 6.2 Main Results
(4) Dialect-Specific Performance Variations. Model effectiveness exhibits notable dependency on the specific SQL dialect, as observed within the Bird-Critic-Multi. Specifically, Gemini-2.0-Flash-Thinking demonstrates the lower performance on PostgreSQL with a 28.99% SR. In contrast, it becomes the most proficient for SQL Server (37.76% SR), with a clear margin over other evaluated models in that dialect. Such variations are plausibly attributable to differential distributions of SQL dialects within the respective training corpora of these models, suggesting that the composition of training data significantly influences dialect-specific debugging capabilities.
@@ 6.4 Trajectory Sampling Comparison
As shown in Table 4, f-Plan Boosting yields $73.7\%$ more successful trajectories than the baseline while maintaining similar runtime and overhead. By contrast, rejection sampling increases success rates modestly but at the cost of $4.2\times$ more attempts and $4.5\times$ longer execution time. When combined, rejection sampling and f-Plan achieve the best overall trade-off, generating 2,560 trajectories with reduced average attempts ($1.7$) and a $62\%$ reduction in cost relative to rejection sampling alone. These results demonstrate that f-Plan provides an effective and efficient approach to trajectory augmentation during rollout in complex environments. Other detailed comparison can be found in Appendix 14.3.
@@ 9 Acknowledgments
10 Environment Setup Details
10.1 SQL Dialects Implementation
@@
10.2 Databases Migration & Modifications
@@
10.3 Issue Collection Protocol
@@
Annotators meticulously review the reproduced issue $(\mathcal{P},\mathcal{S},\sigma_{\text{issue}})$ and craft a new $\sigma^{*}$. This annotation requires ensuring that $\sigma^{*}$: (1) Correctly Implements Intent: Accurately fulfills the user’s objective as inferred from $\mathcal{P}$ and the context of $\sigma_{\text{issue}}$. (2) Resolves the Error: Explicitly fixes the identified flaw(s) in $\sigma_{\text{issue}}$. (3) Is Functionally Correct: Executes successfully on the target database instance $D$ (conforming to $\mathcal{S}$) within the specified dialect and produces the expected, correct results. (4) Adheres to Best Practices: Solution SQLs should present a reasonably efficient and well-formed query. As shown in Figure 1, this results in a curated "Solution Query" ($\sigma^{*}$) paired with the user query and issue SQLs. Finally, to ensure robust evaluation, we annotate each task with evaluation scripts consisting of specific test cases written by Python and SQLs. Details can be found in Appendix 12.
11 Annotator Qualification
11.1 Annotator Entrance Test
@@ 11.3 Qualification Test
12 Evaluation Script Details
@@ 13 Evaluation Metrics
14 More Statistics
14.1 Comparison of BIRD-CRITIC with other conversational Text-to-SQL benchmarks
@@
14.2 Detailed Statistics of BIRD-CRITIC-Multi
@@ 14.3 Quality Validation of Six-Gym
15 Error Analysis Details
@@
16 Experiment Details
16.1 Alias of LLMs
@@
16.2 Model Implementation Details
@@
16.3 Agent Implementation Details
@@
17 Algorithm
17.1 SQL Rewind Algorithm
@@
17.2 Bird-Fixer Algorithm
@@
18 Limitation And Future Work
@@ 20 Prompt
21 BIRD-Fixer Example

### v4

@@ 3 BIRD-CRITIC Benchmark
BIRD-CRITIC is developed via a multi-stage annotation converting raw user issues into executable, verifiable tasks. This involves two annotation groups: 1) 10 qualified database/SQL annotators, who pass strict entry test as detailed in Appendix B.1 and systematic training shown in Appendix B.2 to promise the quality of annotation; 2) 3 senior database experts/scientists for final data collection decision. This process is visually outlined in Figure 2.
@@
We leverage relational databases from the BIRD-SQL development set [24] chosen for its domain diversity across real data-science tasks (California Schools, Financial, Superheroes) and its permissive license. We migrate their original SQLite schemas to PostgreSQL, MySQL, SQL Server, and Oracle, four widely used production-grade dialects. During migration, we go beyond direct dialect translation by refining table and column names. We adjust data types and introduce guarded alterations to schema components to reduce potential information leakage (see Appendix A.2). To pair these databases with realistic debugging scenarios, we collect SQL issue queries from Stack Overflow, following a strict protocol shown in Appendix A.3.
@@
Annotators carefully review the reproduced issue $(\mathcal{P},\mathcal{S},\sigma_{\text{issue}})$ and craft a new $\sigma^{*}$. This annotation requires ensuring that $\sigma^{*}$ can accurately fulfill the user’s objective as inferred from $\mathcal{P}$ and the context of $\sigma_{\text{issue}}$. Also, to ensure robust evaluation, each task is annotated with evaluation scripts consisting of specific test cases written by Python and SQLs. Details can be found in Appendix C. We report the Success Rate (SR %), considering a task solved only when $\sigma_{\text{pred}}$ successfully passes all test cases in its evaluation script.
@@
Figure 3: Distribution of issue categories in all BIRD-CRITIC, derived from an analysis of SQL usage in the real-world database applications. A detailed distribution is in Appendix E.
@@
Table 3 summarizes the key properties of the BIRD-CRITIC benchmark, and Figure 3 visualizes the distribution of its underlying knowledge categories. The distribution of benchmark, is detailed in Appendix E. A side-by-side comparison with standard text-to-SQL benchmarks (Table 6 in Appendix E.1) exposes three distinctive challenges introduced by BIRD-CRITIC: non-query-like problems, multi-dialect complexities, and the most verbose but authentic user queries. As far as we know, BIRD-CRITIC is the first debugging benchmark for SQL applications. These aspects establish BIRD-CRITIC as a crucial benchmark for rigorously evaluating LLM proficiency in solving authentic SQL issues.
@@ 4 Six-Gym: An Automated SQL Debugging Environment for LLMs
GYM-like datasets have proven effective for training LLMs as agents for complex tasks [28]. However, manually collecting and annotating these datasets is labor-intensive and difficult to scale, especially for debugging tasks. Thus, we introduce SQL-Rewind, which addresses this by inverting the debugging paradigm: starting with correct SQL queries ($\sigma^{*}$) and systematically introducing realistic issues to generate issue SQLs ($\sigma_{\text{issue}}$) and user issue query $\mathcal{P}$. This approach enables efficient creation of large-scale training data without human annotation. The pseudo-algorithm is shown in Appendix H.1.
@@ 6.1 SetUp
We evaluate the performance of several popular and strong LLMs across two primary categories, including general-purpose models: Gemini-2.0-Flash, GPT-4.1, Claude-3.7-Sonnet, Qwen-2.5-Coder-32B, Meta-Llama-3.1-8B, Meta-Llama-3.3-70B, Phi-4 and DeepSeek-V3. The second category consists of models specifically renowned for their advanced reasoning capabilities: O3-mini, O1-preview, Gemini-2.0-Flash-Thinking, and Claude-3.7-Sonnet-Thinking. The implementation details are in Appendix G.2.
@@ 6.2 Main Results
(4) Dialect-Specific Performance Variations. Model effectiveness exhibits notable dependency on the specific SQL dialect, as observed within the Bird-Critic-Multi. Specifically, Gemini-2.0-Flash-Thinking demonstrates the lower performance on PostgreSQL with a 28.99% SR. In contrast, it becomes the most proficient for SQL Server (37.76% SR), with a clear margin over other evaluated models in that dialect. Such variations are plausibly attributable to differential distributions of SQL dialects within the respective training corpora of these models, suggesting that the composition of training data significantly influences dialect-specific debugging capabilities.w
@@ 6.4 Trajectory Sampling Comparison
As shown in Table 4, f-Plan Boosting yields $73.7\%$ more successful trajectories than the baseline while maintaining similar runtime and overhead. By contrast, rejection sampling increases success rates modestly but at the cost of $4.2\times$ more attempts and $4.5\times$ longer execution time. When combined, rejection sampling and f-Plan achieve the best overall trade-off, generating 2,560 trajectories with reduced average attempts ($1.7$) and a $62\%$ reduction in cost relative to rejection sampling alone. These results demonstrate that f-Plan provides an effective and efficient approach to trajectory augmentation during rollout in complex environments. Other detailed comparison can be found in Appendix E.3.
@@ 9 Acknowledgments
Appendix A Environment Setup Details
A.1 SQL Dialects Implementation
@@
A.2 Databases Migration & Modifications
@@
A.3 Issue Collection Protocol
@@
Annotators meticulously review the reproduced issue $(\mathcal{P},\mathcal{S},\sigma_{\text{issue}})$ and craft a new $\sigma^{*}$. This annotation requires ensuring that $\sigma^{*}$: (1) Correctly Implements Intent: Accurately fulfills the user’s objective as inferred from $\mathcal{P}$ and the context of $\sigma_{\text{issue}}$. (2) Resolves the Error: Explicitly fixes the identified flaw(s) in $\sigma_{\text{issue}}$. (3) Is Functionally Correct: Executes successfully on the target database instance $D$ (conforming to $\mathcal{S}$) within the specified dialect and produces the expected, correct results. (4) Adheres to Best Practices: Solution SQLs should present a reasonably efficient and well-formed query. As shown in Figure 1, this results in a curated "Solution Query" ($\sigma^{*}$) paired with the user query and issue SQLs. Finally, to ensure robust evaluation, we annotate each task with evaluation scripts consisting of specific test cases written by Python and SQLs. Details can be found in Appendix C.
Appendix B Annotator Qualification
B.1 Annotator Entrance Test
@@ B.3 Qualification Test
Appendix C Evaluation Script Details
@@ Appendix D Evaluation Metrics
Appendix E More Statistics
E.1 Comparison of BIRD-CRITIC with other conversational Text-to-SQL benchmarks
@@
E.2 Detailed Statistics of BIRD-CRITIC-Multi
@@ E.3 Quality Validation of Six-Gym
Appendix F Error Analysis Details
@@
Appendix G Experiment Details
G.1 Alias of LLMs
@@
G.2 Model Implementation Details
@@
G.3 Agent Implementation Details
@@
Appendix H Algorithm
H.1 SQL Rewind Algorithm
@@
H.2 Bird-Fixer Algorithm
@@
Appendix I Limitation And Future Work
@@
Appendix L BIRD-Fixer Example